A drug needs to dissolve a 300.0-g solid substance in a 650-ml liquid substance (ρ = 0.87 g/ml) before it could be administered. What could be the density of the resulting solution if there were no changes in volume after the solid substance was dissolved?
1
Expert's answer
2020-10-02T07:14:22-0400

Mass of the liquid subctance "m=650\cdot 0.87=565.5(g)"


The density of the resulting solution "(V=650 ml=const)"


"\rho=(300+565.5)/650=1.33(g/ml)"